OpenStudy (anonymous):

Which of the following numbers could be used as a common denominator for adding the fractions 2/63,-4/35,8/21

OpenStudy (cwrw238):

fid prime factors of the 3 deomiatots 63 = 3*3*7 21 = 3*7 35 = 5 * 7 7 is commo to all 3 ad is oly used oce LCM = 3*3*3*5*7 = 27 * 35

OpenStudy (cwrw238):

- its to that i cat spell - the letter after m o my keyboard is to working lol

OpenStudy (anonymous):

aha its okayy thank you